# Unstructured Evidence Attribution for Long Context Query Focused Summarization

### v1

@@ 4 Experiments and Results
We use four test datasets (full dataset descriptions in Appendix B). At a high level these are: SQuALITY (Wang et al. 2022, short sci-fi novels, single document, average context length: 5,200 tokens); LexAbSumm (Santosh et al. 2024, long legal documents, single document, average context length: 14,357 tokens); SummHay (Laban et al. 2024, synthetic conversations and news, multi-document, average haystack context length: 93,000 tokens); and ScholarQABench (Asai et al. 2024, Computer Science research papers, multi-document, average context length: 16,341 tokens).
@@
We use a set of LLMs covering multiple sizes and pretraining configurations. This includes Llama 3.2 1B, Llama 3.2 3B, Llama 3.1 8B Dubey et al. (2024), Mistral Nemo 2407, and Mixtral 8x7B. [^2] Huggingface model IDs are listed in Appendix D Table 5 Additionally, we provide an upper bound estimate on performance using GPT 4o mini with no fine-tuning.
(a) Llama 3.2 1B
(b) Llama 3.2 3B
(c) Llama 3.1 8B
@@
(f) GPT 4o Mini
Figure 4: Location of extracted evidence in the provided source context for different methods.
(a) SQuALITY
(b) LexAbSumm
(c) SummHay
(d) ScholarQABench
Figure 5: Location of ground truth evidence in each dataset.
SLTS LASS SMHM SQBM
Method Rel Con Rel Con Rel Con Rel Con
Llama 3.2 1B 2.68 2.15 3.68 3.38 4.53 4.40 3.80 3.61
+ Standard 2.73= 2.17= 3.25 2.93 4.53= 4.44= 3.81= 3.59=
+ Shuffled 2.79* 2.15= 3.41 3.03 4.66* 4.55* 3.97* 3.69*
Llama 3.2 3B 4.39 4.05 4.40 4.19 4.82 4.74 4.28 4.11
+ Standard 4.22 3.80 4.19 4.02 4.90* 4.85* 4.41* 4.21*
+ Shuffled 3.84 3.38 4.25 4.02 4.89* 4.84* 4.49* 4.23*
Llama 3.1 8B 4.55 4.34 4.64 4.52 4.88 4.78 4.18 4.06
+ Standard 4.63* 4.41* 4.53 4.44 4.94* 4.93* 4.64* 4.42*
+ Shuffled 4.59* 4.34= 4.55 4.44 4.97* 4.92* 4.68* 4.41*
Mistral Nemo 2407 4.27 4.09 3.83 3.85 4.27 4.15 3.15 3.23
+ Standard 4.43* 4.24* 4.03* 4.04* 4.54* 4.47* 3.79* 3.75*
+ Shuffled 4.53* 4.35* 4.18* 4.12* 4.65* 4.49* 3.49* 3.41*
Mixtral 8x7B 4.02 3.99 4.28 4.22 4.78 4.68 3.95 3.89
+ Standard 4.52* 4.35* 4.45* 4.40* 4.84* 4.72* 4.26* 4.13*
+ Shuffled 4.51* 4.40* 4.44* 4.38* 4.79= 4.68= 4.33* 4.18*
GPT 4o Mini 4.98 4.92 4.93 4.77 4.99 4.98 4.94 4.76
Table 4: Relevance and consistency of generated summaries. Relevance and consistency are measured using an autorater (GPT-4o-mini) Liu et al. (2023) based on previously validated prompts Liu et al. (2024b). * indicates significance as measured by non-overlapping bootstrapped confidence intervals with the baseline. = indicates no significant difference from baseline. S indicates single document tasks, M indicates multi-document. SLT is SQuALITY, LAS is LexAbSumm, SH is SummHay, and SQB is ScholarQABench.
Evaluation
For evaluation, we follow recent trends in summarization evaluation, which have noted that traditional lexical based metrics such as ROUGE score Lin (2004) are insufficient for more complex summarization tasks Koh et al. (2022). We evaluate our models using autoraters (i.e., LLM-as-a-judge) Gu et al. (2024); Zheng et al. (2023); Liu et al. (2023) along two dimensions using previously validated prompts listed in Appendix A (Figure 20 and Figure 21Liu et al. (2024b). These dimensions are Relevance and Consistency. Given a source text, a target text, and optionally a query, Relevance measures how well the target covers the main points of the source, as well as how much irrelevant or redundant information it contains. Consistency measures to what degree the target contains any factual errors with respect to the source. Both scores are measured on a scale from 1-5 using GPT-4o-mini OpenAI (2023). [^3] We test the robustness of our evaluation in Appendix F
Training, and Inference
@@
4.1 RQ1: Can LLMs Generate Unstructured Evidence?
Using the datasets and models just described, we first test if LLMs can extract and effectively utilize unstructured evidence, as well as the impact of training on SUnsET. We look at two aspects of citation ability: evidence hallucination and evidence accuracy.
To study evidence hallucination, we attempt to match each sentence generated in the evidence list to its position in the context. We do so with two measures: exact string match and percent overlap of longest common substring (LCS) between the evidence and context. We present the rate of exact evidence match and 50% LCS overlap for all models aggregated across all datasets in Table 2. We see that all base models struggle to faithfully copy evidence from the context. This includes GPT 4o mini, which only faithfully copies 11% of the time. This rate is dramatically improved in all cases except for the smallest model (Llama 3.2 1B) by learning to cite unstructured evidence using SUnsET. Additionally, we see that rates of citation also dramatically increase (6.8$\times$ for Mixtral 8x7B). We find that learning to cite using SUnsET greatly improves the extraction of unstructured evidence from arbitrary contexts.
[…]
We present results on citation relevance and consistency in Table 3. We again find that without any intervention, base models are generally bad at selecting and generating relevant and consistent evidence. As expected, larger models are better at this task, with base GPT 4o mini providing a modest upper bound on performance. We see that training on SUnsET helps to significantly close this gap and greatly improve citation quality. Smaller models see fewer gains, while larger models are able to adapt using SUnsET much more strongly, in some cases surpassing GPT 4o mini (e.g., Llama 3.1 8B for ScholarQABench). As with evidence hallucination, standard fine-tuning generally performs better, except for Mixtral, which sees a boost from doing shuffled training. Overall, we find that base models struggle to utilize unstructured evidence, while SUnsET helps models to learn this across highly diverse test sets.
(a) Llama 3.2 1B
(b) Llama 3.1 8B
(c) Mixtral 8x7B
(a) Llama 3.2 1B
(b) Llama 3.1 8B
(c) Mixtral 8x7B
@@ 4.2 RQ2: Is evidence lost-in-the-middle?
Next, we explore to what extent this evidence is lost in the middle. To characterize positional bias, we match extracted evidence to its relative location in the document context (based on 50% LCS overlap) and plot this as a histogram in Figure 4. As a point of reference, we also plot the distribution of summary sentence locations within the test set documents by matching ground truth reference summaries to their relative locations in their context documents in Figure 5.444We find the relative location using cosine similarity of S-BERT sentence embeddings Reimers and Gurevych (2022)
We find that evidence is lost in the middle for all base models. This includes GPT 4o Mini, which has a sharp spike of evidence in the early context. This stands in contrast to ground truth summary location distributions, which are uniform in all cases except for LexAbSumm which has a bias for evidence at the end of the context. In general, training on SUnsET without shuffling increases the rate of evidence extraction, but does not decrease the bias significantly. Shuffling on the other hand, increases the rate of evidence extraction and can decrease the bias. This is especially the case for Llama 3.1 8B and Llama 3.2 3B. Thus, similar to previous work on RAG for question-answering tasks Zhang et al. (2024b), we find that shuffled training has the potential to help reduce positional biases for evidence extraction as well. This presents a tradeoff between training with shuffled and unshuffled documents: on the one hand, standard training leads to generally higher intrinsic citation and evidence quality, while on the other, fails to reduce positional bias. Shuffling reduces positional bias, potentially utilizing more relevant evidence for the final summary, but suffers a penalty in terms of citation and evidence quality.
@@ 4.3 RQ3: Is Summary Quality Improved?
Finally, we test if learning to cite has a positive impact on summary quality. For this experiment, we measure the relevance and consistency of every summary with respect to its context and query. We again compare each base model to training on SUnsET with standard and shuffled context. Our results are presented in Table 4.
We find that summary quality is uniformly and significantly improved by learning to cite unstructured evidence. Larger models adapt more easily than smaller models which struggle on the single-document datasets (SQuALITY and LexAbSumm). Learning from SUnsET has an especially strong impact on multi-document datasets, while standard and shuffled training generally lead to similar gains in performance. Mistral Nemo tends to perform better with shuffled training, while for other models the results are mostly interchangeable. The selection of which approach is best then comes down to the tradeoff between evidence quality and positional bias. Ultimately though, fine-tuning on SUnsET helps reduce the performance gap in terms of summary quality with much more powerful models such as GPT 4o Mini.
Additionally, for LexAbSumm, we see a drop in performance for Llama models but gains in performance or Mixtral. Recall that LexAbSumm expresses a bias in terms of where relevant information in the summary lies. With the exception of Mixtral, shuffled training tends to mitigate this drop slightly, or lead to the best performance in the case of Nemo, which may be attributed to the reduced evidence bias of this setting.
Finally, To observe the impact of number of SUnsET training samples on summary quality, we plot relevance and consistency vs. number of training samples for SQuALITY and ScholarQABench in Figure 6 and Figure 7. Interestingly, we find that performance generally peaks with only a modest amount of data (around 1k-3k samples depending on the model) at which point performance plateaus or slightly drops. Given this, we see that adapting models to our task requires minimal data and can be performed relatively cheaply.
@@ 5 Discussion and Conclusion
Generating unstructured evidence citations alongside query-focused summaries has the potential to improve user trust and transparency in LLMs. Our study has highlighted salient challenges in this task, as well as a potential solution for them. With no intervention, these models suffer from the lost-in-the-middle problem, which we demonstrate across many settings for the case of unstructured evidence citation. They additionally struggle to generate accurate unstructured evidence from their contexts. Our proposed dataset, SUnsET, serves as a useful domain-agnostic synthetic dataset to help mitigate these issues. This intervention is at training time, meaning the inference cost is lower than for complex reasoning and inference chains. In addition to improving evidence quality, overall summary quality is improved. We hope this work can be built upon to help create more reliable, transparent, and useful summarization systems.
@@ Acknowledgements
While our approach offers several benefits, there are notable areas to improve upon. Generating unstructured evidence directly can be prone to hallucination, while it is critical for the evidence to be exactly correct. A more precise RAG approach may offer some benefits. While shuffling during training helps the model to pull evidence more evenly, this also reduces the benefits in terms of evidence quality. A more targeted approach based on directly altering positional embeddings may be more appropriate for this Hsieh et al. (2024). We experiment with documents using a fixed number of sections in this study; allowing for variable-length documents could deliver greater improvements in performance. Additionally, we acknowledge potential prompt bias, influencing model outputs. Despite our efforts to mitigate these effects, they persist as a challenge, and using techniques such as APO Pryzant et al. (2023) could address these issues. Finally, while SUnsET data is domain agnostic, it could be worth exploring how domain-aware data could help for more targeted applications (e.g., in the legal domain).
@@ A.1 Synthetic Data Generation Prompts
Please write a summary which addresses the question. Please make the summary as specific and detail oriented as possible. Please include actual examples from the book when possible. Please do not write more than is absolutely necessary. After you write the summary, please write exact quotes and passages you will include in the book, from which the summary could be written. Please include at least {n_evidence} of these passages, which you intend to include verbatim in the book. Please indicate the exact chapter where the passages will be written in a separate field. **OUTPUT FORMAT** Please a JSON object with two fields: “summary”, “evidence”, and “chapter”. The summary field should have the summary. The evidence field should have a list of evidence sentences from the book. The chapter field should have the exact chapter where the corresponding evidence sentence will appear. Please only indicate the chapter number for this field. There should be the same number of elements in the “evidence” field as there are in the “chapter” field. In other words, as: ```python { ‘summary’: ‘Summary text’, ‘evidence’: [[[[‘evidence sentence 1’, ‘evidence sentence 2’, …]]]] ‘chapter’: [[[[1, 4, …]]]] } ```
@@
Please add citations to all citation-worthy statements in the essay using the numbered evidence list, by indicating the citation numbers of the corresponding evidence. More specifically, add the citation number at the end of each relevant sentence in the essay before the punctuation mark e.g., ‘This work shows the effectiveness of problem X [[[[1]]]].’ when the passage [[[[1]]]] in the evidence list provides full support for the statement. Only add a citation if it is fully relevant and unambiguously supportive of that sentence. Not all evidences may be relevant, so only cite those that directly support the statement. Please do not add any explanations or justifications for the evidence, simply indicate the evidence numbers if they are relevant. If a sentence does not use any of the provided evidence, please simply copy the sentence as is and do not add anything to the end of it. If multiple evidences support a statement, please cite them together (e.g., [[[[1]]]][[[[2]]]]). For each citation-worthy statement, you only need to add at least one citation, so if multiple evidences support the statement, just add the most relevant citation to the sentence.
@@ A.3 Evaluation Prompts
Your task is to read a document and then write an essay which addresses the following question: {question_text} To write your essay, you should read the document and identify key passages which will help guide your response. Extract every passage which is directly relevant for your essay. Please copy each extracted passage to a list in the format specified below. Please copy the exact text of each passage (do NOT paraphrase!). Then, write your essay which addresses the query. Please add citations to all citation-worthy statements using the extracted evidence, by indicating the citation numbers of the corresponding evidence. More specifically, add the citation number at the end of each relevant sentence before the punctuation mark e.g., ‘This work shows the effectiveness of problem X [1].’ when the passage [1] in the evidence list provides full support for the statement. Only add a citation if it is fully relevant and unambiguously supportive of that sentence. Not all evidences may be relevant, so only cite those that directly support the statement. Please do not add any explanations or justifications for the evidence, simply indicate the evidence numbers if they are relevant. If a sentence does not use any of the provided evidence, please simply copy the sentence as is and do not add anything to the end of it. If multiple evidences support a statement, please cite them together (e.g., [1][2]). For each citation-worthy statement, you only need to add at least one citation, so if multiple evidences support the statement, just add the most relevant citation to the sentence. Please limit to only 10 pieces of evidence. Here is the document: {context} **OUTPUT FORMAT** Output your response as: EVIDENCE: [[[[1]]]] Extracted passage 1 [[[[2]]]] Extracted passage 2 [[[[N]]]] Extracted passage N RESPONSE: response
@@
Figure 20: Relevance evaluation prompt from Liu et al. (2024b). The {document} field is filled with the document context and the {summary} field is filled with a summary. When used to evaluate summarization, the {query} field is filled with the query used to generate the summary. For citation evaluation, the {query} field and all references to queries are removed from the prompt.
@@
Figure 21: Consistency evaluation prompt from Liu et al. (2024b). The {document} field is filled with the document context and the {summary} field is filled with a summary. When used to evaluate summarization, the {query} field is filled with the query used to generate the summary. For citation evaluation, the {query} field and all references to queries are removed from the prompt.

### v2

@@ 4 Experiments and Results
Figure 5: Average relevance and consistency of evidence texts with respect to their citation sentences measured using an autorater (DeepSeek-V3; Liu et al., 2023) based on prompts which have previously undergone human evaluation for quality Liu et al. (2025). Bold indicates best performance for a given model; “*” and “+” indicate statistical significance above the fixed granularity and non-fine-tuned unstructured baselines, respectively, based on non-overlapping 95% confidence intervals.
@@
We use four test datasets (full descriptions in Appendix B). These include three human written datasets, forcing models trained on SUnsET to generalize beyond synthetic data. These are: SQuALITY (Wang et al. 2022, short sci-fi novels, single document, average context length: 5,200 tokens); LexAbSumm (Santosh et al. 2024, long legal documents, single document, average context length: 14,357 tokens); SummHay (Laban et al. 2024, synthetic conversations and news, multi-document, average haystack context length: 93,000 tokens); and ScholarQABench (Asai et al. 2024, Computer Science research papers, multi-document, average context length: 16,341 tokens). We present here the average results from sampling evenly across datasets, results on individual datasets are presented in Appendix D.
@@
We test Llama 3.2 1B, Llama 3.2 3B, Llama 3.1 8B Dubey et al. (2024), Mistral Nemo 2407, and Mixtral 8x7B. [^2] Huggingface model IDs are listed in Appendix G Table 8 We compare four settings for each LLM: base models with fixed granularity evidence (Fixed Gran.), base models with unstructured evidence citation (Unstruct. Base), training adapters on SUnsET (+ SunSET), and training adapters on shuffled SUnsET documents (+ Shuffled). Additionally, we provide an upper bound estimate on performance using GPT 4o mini with no fine-tuning.
Model Exact Match 50% Match # Evidence
Llama 3.2 1B  0.0 35.71     14
+ SUnsET   7.69 43.26   208
+ Shuffle   5.15 22.68     97
Llama 3.2 3B 25.57 90.11 1345
+ SUnsET 52.77 85.62 3720
+ Shuffle 32.99 74.07 2337
Llama 3.1 8B 43.93 83.12 3412
+ SUnsET 78.36 97.21 4690
+ Shuffle 54.53 88.51 4684
Mistral Nemo 2407   5.48 66.13   310
+ SUnsET 82.20 97.29 2107
+ Shuffle 72.38 95.76 1959
Mixtral 8x7B   5.79 91.25 3452
+ SUnsET 33.82 90.47 4208
+ Shuffle 29.29 90.74 4288
GPT-4o-mini 11.06 96.32 8159
Table 3: Evidence copy rates. We measure exact string match (i.e. when the evidence sentence exactly appears in the context) as well as 50% overlap between the extracted evidence and the longest common substring.
Evaluation
We evaluate our models using autoraters Gu et al. (2024); Zheng et al. (2023); Liu et al. (2023) along two dimensions. These dimensions are Relevance and Consistency. Given a source text, a target text, and optionally a query, Relevance measures how well the target covers the main points of the source, as well as how much irrelevant or redundant information it contains. Consistency measures to what degree the target contains any factual errors with respect to the source. Both scores are measured on a scale from 1-5 using DeepSeek-V3 Liu et al. (2024a). [^3] We validate the robustness of the ratings from DeepSeek-V3 in Appendix I. We use prompts which have been previously validated to correlate well with human annotations of relevance and consistancy (listed in Appendix A Figure 21 and Figure 22Liu et al. (2025).
Figure 6: Relevance and consistency F1 scores. Bold best performance for a given model; “*” and “+” indicate statistical significance above the fixed granularity and non-fine-tuned unstructured baselines, respectively, based on non-overlapping 95% confidence intervals.
(a) Llama 3.2 3B
(b) Llama 3.1 8B
(c) GPT 4o Mini
@@
(f) Test Datasets
Figure 7: Distribution of location of extracted evidence in the provided source context for different methods. Test dataset evidence location is measured by comparing to reference summaries.
Figure 8: Relevance and consistency of generated summaries. Bold best performance for a given model; “*” and “+” indicate statistical significance above the fixed granularity and non-fine-tuned unstructured baselines, respectively, based on non-overlapping 95% confidence intervals.
4.1 RQ1: Can LLMs Use Unstructured Evidence?
Using the datasets and models just described, we first test how well models can copy and utilize unstructued evidence (i.e., any span of arbitrary length from the context). We look at two aspects: evidence copy accuracy, and evidence quality.
Copy Accuracy
To study copy accuracy, we match each piece of evidence to its longest common substring (LCS) in the context. We present the rate of exact evidence match and 50% LCS overlap for all models aggregated across all datasets in Table 3. We see that without fine-tuning, models struggle to copy evidence from the context. This includes GPT 4o mini, which only copies perfectly 11% of the time. SUnsET helps models learn to copy evidence spans in all cases except for the smallest model (Llama 3.2 1B). We see that the number of citations also dramatically increases.
Evidence Quality
@@ 4.1 RQ1: Can LLMs Use Unstructured Evidence?
We find that the average citation quality of unstructured evidence is better than fixed granularity evidence (Figure 5). This validates the unstructured evidence approach, where flexible evidence extraction enables higher quality citations to source texts. We also see that models’ ability to extract quality evidence is improved by SUnsET, where our results are on par with GPT 4o Mini. When balancing citation quality and citation quantity (Figure 6), we see that learning to use unstructured evidence with SUnsET leads to statistically significant improvements over fixed-granularity and non-fine-tuned baselines across models. This is particularly the case for medium to larger models. For smaller models (particularly, Llama 3.1 1B), simply fine-tuning for such a complex task is insufficient, where all settings struggle to extract and use evidence. Non-shuffled training is often better than shuffled training, though shuffled training also improves citation quality by a large margin. When balancing for recall, fixed-granularity evidence tends to be better than unstructured evidence without fine-tuning, which makes sense as a model only needs to generate references in the fixed-granularity case. Thus, the primary benefits to citation quality by learning from SUnsET are two-fold: the quality of the evidence itself improves, and the rate of citation improves.
@@ 4.2 RQ2: Is evidence lost-in-the-middle?
Next, we quantify to what extent unstructured evidence is lost in the middle. For this, we match extracted evidence to its relative location in the document context (based on 50% LCS overlap) and plot the distributions in Figure 7. As a point of reference, we also plot the distribution of summary sentence locations within the test set documents by matching ground truth reference summaries to their relative locations in their context documents. [^4] We find the relative location using cosine similarity of S-BERT sentence embeddings Reimers and Gurevych (2019)
We find that evidence is lost in the middle for all non-fine-tuned models, most often appearing at the beginning or end of the context. This includes GPT 4o Mini, which has a sharp spike of evidence in the early context. This stands in contrast to ground truth summary location distributions, which are uniform in all cases except for LexAbSumm which has a bias for evidence at the end of the context. In general, training on SUnsET without shuffling increases the rate of evidence extraction, and can help decrease the bias. Shuffling on the other hand, increases the rate of evidence extraction and decreases the bias in all cases except for Mixtral 8x7B. Thus, the modular nature of SUnsET documents, where global structure can be shuffled while local structure is maintained, can be utilized to help reduce positional biases in evidence selection, better reflecting the natural distribution of evidence based on reference data.
@@ 4.3 RQ3: Is Summary Quality Improved?
Finally, we test if using unstructured evidence has a positive impact on summary quality. To do so, we measure the relevance and consistency of every summary with respect to its context and query. Our results are presented in Figure 8 (results on individual datasets are given in Appendix D).
First, for fixed granularity evidence the summaries tend to be similar or slightly lower in quality than unstructured with no fine-tuning, further motivating the unstructured approach. This is likely because the unstructured evidence task has two subtasks: salient evidence selection, followed by summarization, which has been linked to improvements in summary quality Ernst et al. (2024). Second, we find that training on SUnsET leads to statistically significant improvements in summary quality over both baselines. Standard and shuffled training on SUnsET generally lead to similar gains in performance over unstructured with no fine-tuning, meaning the selection of which approach comes down to a tradeoff between overall evidence quality (where standard has a slight edge) and evidence diversity (where shuffled has an edge). To observe the effect of number of training samples from SUnsET, we perform an ablation where we fine-tune on different number of samples in Appendix E Figure 23 and Figure 24, finding that best performance only requires around 3k samples. Third, by measuring Pearson’s R correlation between citation and summary scores, we find a moderate correlation (0.35 for Relevance and 0.34 for Consistency), demonstrating a relationship between the quality of the citations and the quality of the summaries. Ultimately, we show the unstructured evidence setup can lead to better evidence and summaries, and demonstrate the utility of SUnsET for learning the task across diverse, human written data.
@@ 5 Discussion and Conclusion
Citing precise evidence spans of any arbitrary length for LCQFS has the potential to improve user trust in LLM summaries, as well as the quality of the evidence. Our study highlights salient challenges in this task, contrasts it with the fixed-granular approach, and demonstrates an effective method towards solving it. With no intervention, evidence is lost-in-the-middle, which we show across many settings for the case of unstructured evidence. They additionally struggle to accurately copy arbitrary length evidence from their contexts by default. Our proposed dataset, SUnsET, serves as a useful and inexpensive synthetic dataset to mitigate these issues. This intervention is at training time, meaning the inference cost is lower than for complex reasoning and inference chains. In addition to improving evidence quality, overall summary quality is improved. We hope this work can be built upon to help create more reliable, trustworthy, and useful summarization systems.
@@ Acknowledgements
While our approach offers several benefits, there are notable areas to improve upon. Generating unstructured evidence directly can be prone to hallucination, while it is critical for the evidence to be exactly correct. A more precise RAG approach may offer some benefits. While shuffling during training helps the model to pull evidence more evenly, this also reduces the benefits in terms of evidence quality. A more targeted approach based on directly altering positional embeddings may be more appropriate for this Hsieh et al. (2024). We experiment with documents using a fixed number of sections in this study; allowing for variable-length documents could deliver greater improvements in performance. Additionally, we acknowledge potential prompt bias influencing model outputs, and that synthetic data may have characteristics which differ from human-written texts. Despite our efforts to mitigate these effects, they persist as a challenge, and using techniques such as APO Pryzant et al. (2023) could address these issues. Finally, while SUnsET data is domain agnostic, it could be worth exploring how domain-aware data could help for more targeted applications (e.g., in the legal domain).
@@ A.1 Synthetic Data Generation Prompts
Please write a summary which addresses the question. Please make the summary as specific and detail oriented as possible. Please include actual examples from the book when possible. Please do not write more than is absolutely necessary. After you write the summary, please write exact quotes and passages you will include in the book, from which the summary could be written. Please include at least {n_evidence} of these passages, which you intend to include verbatim in the book. Please indicate the exact chapter where the passages will be written in a separate field. **OUTPUT FORMAT** Please a JSON object with two fields: “summary”, “evidence”, and “chapter”. The summary field should have the summary. The evidence field should have a list of evidence sentences from the book. The chapter field should have the exact chapter where the corresponding evidence sentence will appear. Please only indicate the chapter number for this field. There should be the same number of elements in the “evidence” field as there are in the “chapter” field. In other words, as: ```python { ‘summary’: ‘Summary text’, ‘evidence’: [[‘evidence sentence 1’, ‘evidence sentence 2’, …]] ‘chapter’: [[1, 4, …]] } ```
@@
Please add citations to all citation-worthy statements in the essay using the numbered evidence list, by indicating the citation numbers of the corresponding evidence. More specifically, add the citation number at the end of each relevant sentence in the essay before the punctuation mark e.g., ‘This work shows the effectiveness of problem X [[1]].’ when the passage [[1]] in the evidence list provides full support for the statement. Only add a citation if it is fully relevant and unambiguously supportive of that sentence. Not all evidences may be relevant, so only cite those that directly support the statement. Please do not add any explanations or justifications for the evidence, simply indicate the evidence numbers if they are relevant. If a sentence does not use any of the provided evidence, please simply copy the sentence as is and do not add anything to the end of it. If multiple evidences support a statement, please cite them together (e.g., [[1]][[2]]). For each citation-worthy statement, you only need to add at least one citation, so if multiple evidences support the statement, just add the most relevant citation to the sentence.
@@ A.3 Evaluation Prompts
Your task is to read a document and then write an essay which addresses the following question: {question_text} To write your essay, you should read the document and identify key passages which will help guide your response. Extract every passage which is directly relevant for your essay. Please copy each extracted passage to a list in the format specified below. Please copy the exact text of each passage (do NOT paraphrase!). Then, write your essay which addresses the query. Please add citations to all citation-worthy statements using the extracted evidence, by indicating the citation numbers of the corresponding evidence. More specifically, add the citation number at the end of each relevant sentence before the punctuation mark e.g., ‘This work shows the effectiveness of problem X [1].’ when the passage [1] in the evidence list provides full support for the statement. Only add a citation if it is fully relevant and unambiguously supportive of that sentence. Not all evidences may be relevant, so only cite those that directly support the statement. Please do not add any explanations or justifications for the evidence, simply indicate the evidence numbers if they are relevant. If a sentence does not use any of the provided evidence, please simply copy the sentence as is and do not add anything to the end of it. If multiple evidences support a statement, please cite them together (e.g., [1][2]). For each citation-worthy statement, you only need to add at least one citation, so if multiple evidences support the statement, just add the most relevant citation to the sentence. Please limit to only 10 pieces of evidence. Here is the document: {context} **OUTPUT FORMAT** Output your response as: EVIDENCE: [[1]] Extracted passage 1 [[2]] Extracted passage 2 [[N]] Extracted passage N RESPONSE: response
@@
Figure 21: Relevance evaluation prompt from Liu et al. (2025). The {document} field is filled with the document context and the {summary} field is filled with a summary. When used to evaluate summarization, the {query} field is filled with the query used to generate the summary. For citation evaluation, the {query} field and all references to queries are removed from the prompt.
@@
Figure 22: Consistency evaluation prompt from Liu et al. (2025). The {document} field is filled with the document context and the {summary} field is filled with a summary. When used to evaluate summarization, the {query} field is filled with the query used to generate the summary. For citation evaluation, the {query} field and all references to queries are removed from the prompt.